OpenStudy (anonymous):

Write an equation for a circle with a diameter of that has endpoints at (8,7) and (-4,-3). Round to the nearest tenth if necessary.

OpenStudy (anonymous):

I think the answer's (x-y)^2+(y-2)^2=244 is this accurate?

terenzreignz (terenzreignz):

What do endpoints mean?

terenzreignz (terenzreignz):

Does it mean that the line segment from (-4, -3) to (8, 7) is a diameter?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well, that's good for us :) the endpoints give us the radius AND the center of the circle Shall I demonstrate, or can you do it from here?

OpenStudy (anonymous):

Well. Am I right or am I wrong? :P

terenzreignz (terenzreignz):

well, actually, no, you have an (x - y) in your equation, maybe you'd like to correct that :P

terenzreignz (terenzreignz):

Unless you did that on purpose, though

OpenStudy (anonymous):

Oops. :P

OpenStudy (anonymous):

Give me a minute, please.

OpenStudy (anonymous):

(x+2)^2+(y+2)^2 = 61 There, is that better? :)

terenzreignz (terenzreignz):

might I ask how you got it?

OpenStudy (anonymous):

I really am bad at articulating stuff like that, sorry.

OpenStudy (anonymous):

My mind just... cannot transfer that into words. Plus, English isn't my native language.

terenzreignz (terenzreignz):

Well, that's ok, but I'd have to say, your equation was wrong... would you like me to show you the steps to the correct one, or would you prefer to do it yourself?

OpenStudy (anonymous):

I'd love to learn how to answer this question properly and accurately, yes. :)

terenzreignz (terenzreignz):

So you have your endpoints... Let's use them to get the diameter of the circle, with our tool called the distance formula :D The distance from (-4, -3) to (8, 7) is given by \[\sqrt{(7+3)^{2} + (8+4)^{2}}\] Did you do this?

OpenStudy (anonymous):

What....

OpenStudy (anonymous):

I don't know how to find the square root of that.

terenzreignz (terenzreignz):

But you do get \[\sqrt{244}\] right?

OpenStudy (anonymous):

Sorry, that's wrong. It's not one of the multiple choice options on my homework

terenzreignz (terenzreignz):

Because that's not the answer, I only asked you if that's what you get after applying the distance formula :P

OpenStudy (anonymous):

Nope. :P This is a homework question, it's graded on participation, can you please just fully explain the process of this problem? THANK YOU! :)

OpenStudy (anonymous):

I'll be honest, i kind of day dremaed in class yesterday. I remember something about y-k squared... something?

terenzreignz (terenzreignz):

So \[\sqrt{244}\] is the diameter, meaning half of that is the radius \[\frac{\sqrt{244}}{2}\] is the radius Let's put that to one side for now. With the endpoints, we can also determine the center of the circle, it is the midpoint of the endpoints. The x and y value of the midpoint is the average of the x and y values of the two endpoints: \[(\frac{8 - 4}{2},\frac{7 - 3}{2}) = (2, 2)\] The equation of a circle is given by \[(x - h)^{2} + (y - k)^{2} = r^{2}\] Where r is the radius and (h,k) are the coordinates of its circle. Plugging in our radius and center, we get \[(x - 2)^{2} + (y - 2)^{2} = (\frac{\sqrt{244}}{2})^{2} \] \[= (x - 2)^{2} + (y - 2)^{2} = \frac{244}{4} = 61\] so the equation is \[(x - 2)^{2} + (y - 2)^{2} = 61\]

OpenStudy (anonymous):

Okay, I'm going to do another one! Can you stay by me and tell if I did it properly?

terenzreignz (terenzreignz):

I live to serve :D LOL go ahead

OpenStudy (anonymous):

Thank you, give me a minute. The end points for this are (7,-4) (1,-10) btw

OpenStudy (anonymous):

Wait... I've gotten to the square root of 160, but I'm stuck after that

terenzreignz (terenzreignz):

Distance between the two endpoints is \[\sqrt{(-10+4)^{2}+(1-7)^{2}}\]

OpenStudy (anonymous):

I've gotten to the square root of 85, but now I'm stuck!

terenzreignz (terenzreignz):

You still get \[\sqrt{160}\] ?

OpenStudy (anonymous):

....... I've gotten to the square of 160, but now I'm stuck!

terenzreignz (terenzreignz):

I don't know how you could possibly get to those square roots... Maybe you're panicking, just relax! :D \[\sqrt{(-10 +4)^{2} + (1-7)^{2}} = \sqrt{(-6)^{2} + (-6)^{2}} = \sqrt{72} = 6\sqrt{2}\]

terenzreignz (terenzreignz):

So now we have our diameter, to get the radius, just divide it by 2 :)

OpenStudy (anonymous):

3 square root of 2

terenzreignz (terenzreignz):

Yes, so we put that one side, for now, and try to figure out the center of the circle...can you do it? :)

OpenStudy (anonymous):

no

terenzreignz (terenzreignz):

The center of the circle is the midpoint of the given endpoints (Makes sense, right? The center of the circle would surely be the center of the diameter as well :D )

terenzreignz (terenzreignz):

So now can you do it?

OpenStudy (anonymous):

no

terenzreignz (terenzreignz):

The x-coordinate of the midpoint is the average of the x-coordinates of the endpoints so the x coordinate is \[\frac{1+7}{2} = 4\] They y coordinate of the midpoint is the average of the y-coordinates of the endpoints, so can you do the y-coordinate? :)